• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Kowalski

    Log entry time 15:58:53 on August 2,2005

    Entry number 150455

    keyword=Shift Summary

    8:00 Left spectrometer back up. Continue production running.
    9:17 MCC beam tests. Target moved to empty.
    10:23 Studies completed. He target back. Beam back.
    10:59 Beam off. MCC troubleshootin.
    11:26 Beam back. Start at 40 micro-A.
    12:01 Discovered left septum was not set up correctly. Changed to the correct current.
    12:01 lowered beam to 38 micro-A. Septum was getting too hot.
    12:40 Hall-C coming on. Lose beam for a while.
    13:10 While waiting for beam, rebooted septum IOC to impliment software change from yesterday.
    13:20 Septum PS rebooted, ramping up to set currents.
    13:50 Septa up to set currents. Still no beam back. Tested software change by making a momentum change. Septa did not move.
    13:51 Beam back. 40 micro-A. Start production running.
    15:16 Septum at 7.67K. Lowered current to 38 micro-A.
    15:46 Start scanner run. Scan finished. No beam trips. Beautiful!!
    16:00 End of shift.